OpenStudy (nali):

In the sequence below, if each term is defined as n^3 - n^2, where n represents the position, what is the missing number? 0, 4, 48, 100, 180

OpenStudy (anonymous):

0, 4, 18, 48, 100, 180

OpenStudy (nali):

can you please show me how u got this answer

OpenStudy (nali):

@myko

OpenStudy (zzr0ck3r):

1^3-1^2 = 0 2^3-2^2 = 4 3^3-3^2= 18

OpenStudy (zzr0ck3r):

sequence is defined by \[n \in \mathbb{N}\] so start at 1 then put in 2 then 3 ...

OpenStudy (nali):

yes thanks i got it